# Does Turbulence at the Correlation Scale Regulate the Statistics of Magnetic Reconnection?

### v1

@@ Abstract
We solve these equations in a 2D box of size $2\pi\times 2\pi$, with equal values of magnetic and fluid Reynolds numbers $R_{\mu}=R_{\nu}=5000$, using a strongly energy-conserving pseudo-spectral code [25], de-aliased using the 2/3 rule, on a real-space grid with resolution $8192\times 8192$. In code units, the wavenumbers range from 1 to $k_{\text{max}}=(2/3)\times 4096$. We ensure that the inequality $k_{\text{max}}/k_{\text{diss}}(t)\geq 3$, is maintained, where $k_{\text{diss}}=R_{\nu}^{\frac{1}{2}}\langle\omega^{2}+j^{2}\rangle^{1/4}$ is the Kolmogorov dissipation wavenumber, and $\langle\cdots\rangle$ denotes averaging over the simulation domain; this condition provides accurate representation of statistics up to at least fourth order [26].
@@
Two, the global turbulence increments are defined as $\Delta\bm{b}(\bm{x},\bm{r})=\bm{b}(\bm{x}+\bm{r})-\bm{b}(\bm{x})$. We define the transverse increment $\Delta{b}_{\perp}(\bm{x},\bm{r})$ and the longitudinal increment $\Delta{b}_{\parallel}(\bm{x},\bm{r})$ as the projection of $\Delta\bm{b}(\bm{x},\bm{r})$ perpendicular and parallel to the lag vector $\bm{r}$, respectively. For simplicity and consistency with the reconnection increments, we will examine the absolute value of the global increments and discard the $\mid\cdots\mid$.” For the present purposes the transverse turbulence increments are examined at two different lags: the average transverse reconnection scale $\lambda_{\text{rec}}$ and the correlation scale $\lambda_{c}$.
@@
If such a scaling is applicable, Eq. (3) implies that the predicted reconnection rates and reconnection temperature change will be about 500 times larger than previous estimates, while heating rates may be 10,000 times larger. Energy dissipation through reconnection therefore is likely to play a major role in dissipating the turbulence energy cascade.
@@
Appendix A Appendix: Determining Reconnection Statistics

### v2

# Factors Controlling the Statistics of Magnetic Reconnection in MHD Turbulence

@@ Abstract
We solve these equations in a 2D periodic box of size $2\pi\times 2\pi$, with equal values of magnetic and fluid Reynolds numbers $R_{\mu}=R_{\nu}=5000$, using a strongly energy-conserving pseudo-spectral code [25], de-aliased using the 2/3 rule, on a real-space grid with resolution $8192\times 8192$. In code units, the wavenumbers range from 1 to $k_{\text{max}}=(2/3)\times 4096$. We ensure that the inequality $k_{\text{max}}/k_{\text{diss}}(t)\geq 3$, is maintained, where $k_{\text{diss}}=R_{\nu}^{\frac{1}{2}}\langle\omega^{2}+j^{2}\rangle^{1/4}$ is the Kolmogorov dissipation wavenumber, and $\langle\cdots\rangle$ denotes averaging over the simulation domain; this condition provides accurate representation of statistics up to at least fourth order [26].
@@
Two, the global turbulence increments are defined as $\Delta\bm{b}(\bm{x},\bm{r})=\bm{b}(\bm{x}+\bm{r})-\bm{b}(\bm{x})$. We define the transverse increment $\Delta{b}_{\perp}(\bm{x},\bm{r})$ and the longitudinal increment $\Delta{b}_{\parallel}(\bm{x},\bm{r})$ as the projection of $\Delta\bm{b}(\bm{x},\bm{r})$ perpendicular and parallel to the lag vector $\bm{r}$, respectively. Note that different from reconnection increments, in this case the lag vector $\bm{r}$ is a constant across the whole simulation. For simplicity and consistency with the reconnection increments, we will examine the absolute value of the global increments and discard the $\mid\cdots\mid$.” For the present purposes the transverse turbulence increments are examined at two different lags: the average transverse reconnection scale $\lambda_{\text{rec}}$ and the correlation scale $\lambda_{c}$.
@@
Data Availability Statement. The data that support the findings of this study are openly available in the Zenodo repository and can be accessed via Ref. [36].
Appendix A Appendix A: Determining Reconnection Statistics
@@
Appendix B Appendix B: Investigating the effects of selection bias
The central result of this paper is that the distribution of magnetic fields responsible for magnetic reconnection in a 2D MHD turbulent system correlates strongly with the global turbulent field at the energy containing scales of the system, as opposed to the turbulent fields at the transverse reconnection scale. The statistical analysis of magnetic reconnection in this letter is based on a total of 622 events. However, the statistical analysis of the global field consists of a significantly larger number of samples compared to the ensemble of reconnection events. This raises the question of whether the observed correlation is a consequence of selection bias. To address this concern, we studied the effect on the PDF of turbulence increments due to random downsampling of the dataset to 622 increments, equal to the total number of reconnection events in our system. In all cases we examined a lag equal to the average transverse reconnection scale. We downsampled the data using two different methods. First, we randomly sampled 622 increments multiple times from the ensemble of global increments at the transverse reconnection scale. Second, we sampled the transverse increments at 622 randomly chosen locations and directions within the system, using a procedure analogous to that used for computing reconnection increments but without restricting the locations to X-points. In Fig 8, we then compared all of these randomly selected increments to the global probability density function (PDF) and found that they closely matched the global PDF. This analysis robustly eliminates the possibility of selection bias.
Figure 8: Probability density function (PDF) of transverse turbulence increments $\Delta b_{\perp}$ at the average transverse reconnection scale $\lambda_{\mathrm{rec}}$. The global PDF (green) is compared with three independent random subsets of 622 samples (orange, blue, brown; first approach) and with 622 transverse increments sampled at random locations within the global system (magenta; second approach). Average values are shown as vertical lines with corresponding colors. Note that the average values are so similar that the vertical lines are nearly indistinguishable from each other.